An observer measures the length of spaceship and finds that it is exactly one-third of its proper length.THe speed of spaceship is:
A: c
B: √2/3c
C: 2/3c
D: 2√2/3c

Grade:12

1 Answers

Khimraj
6 years ago
the length contraction formula is given by
L = Lo\sqrt{1 - c^{2}/v^{2}}
L = L0/3
So
1/3 =\sqrt{1 - c^{2}/v^{2}}
8/9 = c^{2}/v^{2}
So v = (2\sqrt{2})c/3.
Hope it clears. If you like anser then please approve it.